The point at which the tangent line to the curve is x^3 +y^3=a^3 is parallel to y-axis is?

Grade:11

1 Answers

Arun
6 years ago
Dear student
 
find dy/dx
3 x^2 + 3 y^2 dy/dx = 0
hence
dy/dx =  – x^2/y^2
now put this equals to infinity = 1/0 (parallel to y axis)
hence
 
– x^2/y^2 = 1/0
y^2 = 0
y = 0
now 
x^3 = a^3
x = a
 
hence point will be (a, 0)
 
Regards
Arun (askIITians forum expert)